# Current distribution and chirality in the regime of the quantized Hall effect

### v1

@@ III.3 Four-terminal measurements
Figure 8: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements. The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled such that all three theory curves according to Eq. (II.3) with $\mu=\mu_{0}$ are identical following the edge between shaded and white backgrounds.
@@
we plot for each of the three four-terminal measurements shown in Fig. 7 the branch corresponding to the largest current. We slightly scaled the data to compensate for small variations in the contact resistances. After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) and on the plateaus of the QHE. The edge between shaded and white backgrounds corresponds to the prediction of the Drude model for constant $\mu=\mu_{0}$ [dashed black lines in Fig. 7(a)]. Compared to this model prediction, the current is reduced between the plateaus for all three geometries. The reduction is particularly strong for the set-up with contact 6 connected to ground. In the case of chiral transport, contact 6 is the last one to be reached by the current, while most current should flow into contact 3. The reduction of the current into contact 3 goes along with an enhancement of the current into contact 6, cf. Fig. 7(b). This is a clear signature of non-chiral current but, the current reduction into contact 3 (and the current enhancement into contact 6) exceeds even the prediction of our model expressed by Eq. (II.3). Likewise, for the set-up with contact 2 grounded, the reduction of the current into contact 2 is smaller than the reduction predicted by Eq. (II.3). So, a smaller amount of the non-chiral current reaches contact 5 as it is much further away from contact 1 than contact 2.
The details of the measured current distribution clearly depend on the contact configuration. This behavior demonstrates the limits of our model. For the derivation of Eq. (II.3), we assumed an infinitely long Hall bar without additional grounded contacts. Hence, for identical contact resistances, its prediction is independent of the variations between specific configurations of the four-terminal set-up (the resistance of the Hall bar itself at $B=0$ is much smaller than the contact resistances). To overcome the failure of our simplified model, it would be necessary to account for the more complicate steady-state electrostatics of a multi-terminal device, in particular the inhomogeneity of the electric field near a junction. Suitable numerical calculations could clarify this point. For the moment, we conjecture that such a corrected model would support our previous result based on the three-terminal measurement, namely that the Drude model correctly describes the transition regions between plateaus.
IV Screening theory and chiral transport
Figure 9: Coherent and chiral models for the first plateau of the QHE: Sketch of the distribution of the current $I_{xy}$ near $\nu=1$ for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ with $\vec{B}$ pointing into the plane of the 2DES. Color intensities indicate the distribution of the absolute value of the current density into the three branches. Black bent arrows show the direction and relative magnitudes of electron flow related with the charged edges indicated by blue minus signs for $V_{1}<0$. For $V_{1}>0$, current densities remain the same while the edge charges become positive and black arrows indicate the current flow. (a) Prediction of the Landauer Büttiker picture, assuming that current flows in one-dimensional chiral edge channels only. (b) Prediction of the screening theory for the case of edge ICSs (current flows inside ICSs only). (c) Prediction of the screening theory for the case of a bulk ICS. All three cases predict chiral current flow within the 2DES and result in identical currents $I_{1}$, $I_{2}$ and $I_{3}$ into the contacts. In the stem towards contact 1 we sketch the electrochemical potentials of the first Landau level (LL1, solid line) and the related Hall potential ($U_{\text{H}}$, dotted red line).
In this section, we restrict the discussion to the plateau regions of the quantized Hall resistance, for which our multiterminal experiments indicate chiral and coherent transport. The unique characteristics of chiral transport of a 2D conductor is the opposite directions of current flow along the opposite edges of the conductor. The QHE reveals the additional feature of fundamental resistance values being fractions of $h/e^{2}$. These values can be measured with an extremely high accuracy and stability with deviations of less than $3\times 10^{-9}$ having been demonstrated between high resolution measurements of various samples in different laboratories [16]. In order to explain the measured highly accurate resistance values in fractions of $h/e^{2}$, scattering has to be (almost) completely suppressed.
The LBP [1, 2] is a single-particle model based on the assumption that the Landau levels form perfect 1D edge channels, which support chiral transport, while the bulk of the 2DES is assumed to be electrically insulating because of disorder-induced localization. The edge channels are compressible, meaning that there exist occupied and unoccupied states at the chemical potential. The mere assumption of scattering free transport via compressible edge channels as an explanation of the extremely accurate measurements of the fundamental resistance value $h/e^{2}$ is unsatisfactory.
The screening theory offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$, which is based on first principle calculations [7, 8, 9]. Already the earliest purely classical considerations, taking into account the Coulomb interaction, led to the realization of a segmentation of the 2DES confined in a Hall bar in perfectly screened compressible regions and completely unscreened ICSs [3, 4, 5, 6]. Inside the ICSs, scattering is forbidden because the excitation of a carrier costs the energy $\hbar\omega_{\text{c}}$ of the Landau gap. While numerous experiments already support the screening theory [17, 10, 11, 12, 18, 19, 20, 21, 22, 23, 13], here our focus is on the question whether the screening theory is consistent with chiral transport.
In Fig. 9, we sketch the situation for the first plateau ($\nu=1$, $I_{xx}=0$) of the QHE for the case of a three-terminal junction. As a start, in Fig. 9(a), we show the situation assumed in the LBP. All current flows inside the compressible edge channels (solid blue lines). The stem towards contact 1 contains a sketch of the first Landau level (LL1, solid line), which strictly follows the confinement potential for the single particle LBP. The edge channels emerge, where LL1 intersects with the chemical potential, which differs by $V_{\text{H}}$ between the respective edges. The Hall potential $U_{\text{H}}=eV_{\text{H}}$ (dotted line) drops linearly between the edge channels.
To illustrate the screening theory [9], we consider the two limiting cases realized near the ends of the first plateau of the QHE. For simplicity, we neglect disorder effects. Near the lower magnetic field end with $\nu>1$, the screening theory predicts two narrow ICSs roughly resembling Landauer-Büttiker edge channels, cf. Fig. 9(b). Near the high field end of the plateau, where $\nu\leq 1$, the prediction is a single bulk ICS, while the 2DES near the edges remains compressible, cf. Fig. 9(c). Even for the first case with two edge ICSs, in contrast to the LBP, the current flows entirely inside the ICSs, where scattering is suppressed due to the Landau gap of the density of states. In the compressible regions of the 2DES, electrons can scatter and the local electric field is perfectly screened. As a result, the local electric field entirely drops inside the ICSs. The gradient of the Landau levels, i.e., the effective local electric field drives the current according to $\vec{j}=-en_{\text{s}}\,\vec{E}\times\vec{B}/B^{2}$. Therefore, no current flows in the perfectly screened compressible regions. The gradient of the potential energy of the Landau levels equals the gradient of the local electric field. As can be seen in the sketches of the first Landau level in Figures 9b and 9c, the gradient of the Landau levels has opposite directions inside the ICS(s) along the oposite edges of the sample. This is even the case if only a single ICS covers most of the 2DES, cf. Fig. 9(c). Here, $\vec{E}=0$ and $\vec{j}=0$ in the center of the ICS and the current flows in opposite directions on both sides of the ICS. Consequently, even in case of a single bulk ICS, the current flow is still chiral and the local current density is larger near the sample edges, where the confinement potential drops steeply.
V Summary and Outlook
In summary, we have performed multi-terminal current measurements in the regime of the QHE and compared our results with predictions of a coherent and chiral model versus the Drude model. The Drude model perfectly describes our results in the classical regime of small magnetic fields, the coherent and chiral model fits for the regions corresponding to quantized plateaus of the Hall resistance. In the diffusive regions for large magnetic fields, i.e., between quantized plateaus our three-terminal experiment supports the predictions of the Drude model, which includes non-chiral transport. The four-terminal measurements demonstrate the limits of our model but clearly demonstrate non-chiral current contributions away from the plateaus of quantized Hall resistance. This work contributes to a deeper understanding of the microscopic dynamics and the distribution of current density inside realistic mesoscopic devices in the regime of the QHE.

### v2

@@ III.5 Purely experimental confirmation of non-chiral currents between plateaus
Figure 8: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements. The dotted horizontal lines depict the predicted current ratios for filling factors $1\leq\nu\leq 4$. The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq. (II.2) with $\mu=\mu_{0}$ are identical and follow the edge between shaded and white backgrounds. Sketches: Arrows indicate the displayed current branches; colored numbers indicate current carrying and gray numbers floated contacts.
@@
we plot for each of the three four-terminal measurements shown in Fig. 7 only the branch $I_{k}/I_{1}$ corresponding to the largest current. In a chiral model, this is the current flowing into the first grounded contact after emission from contact 1. We slightly scaled the data (based on the values provided in Table 1) to compensate for small variations in the contact resistances. After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) as well as on the plateaus of the QHE. The line linking shaded and white backgrounds corresponds to the prediction of Eq. (II.2) for constant $\mu=\mu_{0}$ [equal to the dashed black lines in Fig. 7(a)].
Compared to the model prediction, between the plateaus the current is reduced for all three geometries, which above we already interpreted in terms of a finite contribution of the non-chiral $I_{xx}$. Because the various contact resistances of our sample are almost identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$, shown in Fig. 8 should be independent of the details of the contacts configuration (as long as the overall number of grounded contacts remains the same). The fact that $I_{k}(B)/I_{1}$ varies between configurations, is a direct experimental prove of non-chiral currents between the quantized plateaus.
This conclusion experts additional support by a closer look at the amounts of current reduction: It systematically depends on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow.
III.6 Formulation of the proof of chiral current flow for quantized Hall resistance
So far, we have proven non-chiral current flow between the quantized plateaus, while for the quantized plateaus we have merely stated that our experimental results are consistent with chiral current flow. However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow.
The proof is based on the coherent and chiral model introduced in Sec. II.1. Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{ij}$ between the contacts. In this model, chiral current flow corresponds to the specific values of conductance coefficients given by Eq. (3). A non-chiral behavior would correspond to different values, including additional non-zero conductance coefficients in Eq. (3). These would inevitably result in different plateau values of the predicted current ratios.
In all our measurements including three or four terminals shown in Fig. 5, Fig. 7 and Fig. 8, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios. This is in particular evident when comparing the horizontal dotted lines in Fig. 5(b) and Fig. 8, which indicate the predictions of Eq. (4) for the corresponding integer filling factors, with the corresponding measured plateau values.
This result stands in contrast to our finding in between plateaus, where the deviations of our measured currents from the prediction of Eq. (II.2) for $\mu B\gg 1$ provide a proof of non-chiral currents.
IV Summary and Outlook
We have performed multiterminal current measurements of a Hall bar in the regime of the QHE. We find that the current flow is chiral in the magnetic field regions of the quantized Hall resistance plateaus but becomes non-chiral in between the plateaus. Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, which offer well understood, unique physical meanings. The classical Drude model serves us as the basis to describe non-chiral diffusive transport, while our second model is based on the generic chiral and coherent transmission problem. The comparison of different contact configurations for more than three terminals allows for a direct experimental proof of non-chiral currents, which confirm our model based results. Our work contributes to the understanding of the current flow inside mesoscopic devices in the regime of the QHE. It thereby confirms the assumption of chiral current flow frequently used for the case of quantized plateaus of the Hall resistance.